OpenStudy (anonymous):

What is the slope of the rose r=2cos2theta at the point theta = pi

OpenStudy (anonymous):

you got the formula for the derivative?

OpenStudy (fibonaccichick666):

Or can you make it into something you can find the derivative of?

OpenStudy (anonymous):

if not we can recreate it fast, since \[y=r\sin(\theta)\] and \[x=r\cos(\theta)\] and \[\frac{dy}{dx}=\frac{\frac{dy}{d\theta}}{\frac{dx}{d\theta}}\]

OpenStudy (anonymous):

I don't have the formula for the derivative.

OpenStudy (anonymous):

product rule \[\frac{dy}{d\theta}=r'\sin(\theta)+r\cos(\theta)\] \[\frac{dx}{d\theta}=r'\cos(\theta)-r\sin(\theta)\]

OpenStudy (anonymous):

and so \[\frac{dy}{dx}\] is the ratio of those two

OpenStudy (anonymous):

Thank you.